princeevee:

What is the interval notation for the compound inequality? −2

princeevee:

@Ultrilliam

dude:

(−2,9] http://prntscr.com/grdfof

princeevee:

can you help with some more dude?

dude:

Sure I can try

princeevee:

1 attachment
dude:

http://prntscr.com/grdhqj

princeevee:

Solve for x. −9<4x+3≤27 Enter your answer, as one inequality,

princeevee:

@dude

dude:

\[-3<x\le \:6\] Do you want work?

princeevee:

no

princeevee:

1 attachment
dude:

Is the answer a single digit or an inequality?

princeevee:

inequality

dude:

\[x<-24\quad \mathrm{or}\quad \:x\ge \:6\]

princeevee:

alright

princeevee:

1 attachment
dude:

Top right option 4dGw5yUVTUqccxDiKVSvGA.png

princeevee:

1 attachment
dude:

http://prntscr.com/grdsfb I think this, I used (2,2) as a point of reference

princeevee:

1 attachment
dude:

Bottom left graph KOXn9SOqSV2-GQRZg_ECYQ.png

princeevee:

1 attachment
dude:

Also bottom left aQgmJSLqQJWU-S8YMmWXtA.png

princeevee:

1 attachment
dude:

First option looks right

princeevee:

1 attachment
princeevee:

its more than one answer

dude:

http://prntscr.com/grdy78

princeevee:

Suzette ran and biked for a total of 34 mi in 3.5 h. Her average running speed was 6 mph and her average biking speed was 12.5 mph. Let x = total hours Suzette ran. Let y = total hours Suzette biked. Use substitution to solve for x and y. Show your work. Check your solution. (a) How many hours did Suzette run? (b) How many hours did she bike?

princeevee:

maybe just help me through this one

dude:

So I would assume you would write \[6x+12.5y=34\] and \[x+y=3.5\] you can use system of equations which you get 1.5 hours ran and 2 hours biked